OpenStudy (anonymous):

messed up on the last one >_< it was supposed to be 7 -x = 21 but with the -x little and up high if that makes any sense

OpenStudy (lalaly):

\[\large{7^{-x}}=21?\]

OpenStudy (anonymous):

yeah it looks like that lalaly

jimthompson5910 (jim_thompson5910):

\[\Large 7^{-x} = 21\] Convert the equation to logarithmic form to get \[\Large \log_{7}(21)=-x\] \[\Large -x=\log_{7}(21)\] Now multiply both sides by -1 to isolate x \[\Large x=-\log_{7}(21)\]

OpenStudy (phi):

if you do this with a calculator, they do not have a log base 7. If you used log (base 10), which calculators do have: log(7^-x)= log(21) -x log(7)= log(21) x = - log(21)/log(7)